(i) Using Cayley Hamilton theorem, find 𝐴




8 βˆ’ 𝐴




7 + 5𝐴




6 βˆ’ 𝐴




5 + 𝐴




4 βˆ’ 𝐴




3 + 6𝐴




2 +




𝐴 βˆ’ 2𝐼 𝑖𝑓 [




1 2 βˆ’2




2 5 βˆ’4




3 7 βˆ’5




]


1
Expert's answer
2022-01-10T13:00:37-0500

Given matrix is "A = \begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2
\end{bmatrix}"


Then according to Cayley Hamilton theorem,

"|A - \lambda I| = 0"


So we will have,


"A - \lambda I = \begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2
\end{bmatrix} - \lambda\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1
\end{bmatrix}"



"|A - \lambda I| = \begin{vmatrix}
 2-\lambda & 1 & 1 \\
 0 & 1-\lambda & 0 \\
 1 & 1 & 2-\lambda
\end{vmatrix} = 0"




Then equation will be

"\lambda^3 - 5\lambda ^2+7\lambda -3=0"

According to Cayley Hamilton theorem,

Every matrix is the root of it's eigen matrix.

then, "A^3 - 5A^2+7A -3=0" (1)


Given equation is "A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5 + A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I"


This equation can be written as,

"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5 + A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I = (A^3 - 5A^2+7A -3)(A^5+A) + (A^2+A+I)"


From equation (1), above equation will be modified as,

"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5 + A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I = A^2+A+I"


So putting value of "A^2,A,I"

we will get,

"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5 + A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I"

"= \begin{bmatrix}
 5 & 4 & 4 \\
 0 & 1 & 0 \\
 4 & 4 & 5
\end{bmatrix} + \begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2
\end{bmatrix} + \begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1
\end{bmatrix} = \begin{bmatrix}8&5&5\\ 0&3&0\\ 5&5&8\end{bmatrix}"